# Learning Actionable World Models for Industrial Process Control

### v1

@@ 4.3 Training details
We train the world model using stochastic gradient descent with batch size $32$ for $500$ epochs using the Adam optimizer [35] with learning rate $\alpha=3\times 10^{-4}$ and the default settings otherwise. For fine-tuning, the model is pre-trained for $10$ epochs. All parameters of $P_{z}$, $P_{a}$ and the last layer of $Enc$ are fine-tuned for $500$ epochs. The coefficients for the loss function in Eq. (4) are set as $\lambda_{1}=1,\lambda_{2}=10$, balancing the terms such that they are of roughly equal magnitude.
5 Experimental setup and evaluation
5.1 Experiments
@@
Experiment 1: Action space coverage. The goal is to learn to what extent the world model can learn the effects of actions across the whole action space, based on a sparse sampling of all possible actions. We conduct this analysis individually on datasets D1 and D2 that offer a systematic coverage of the action space. We use as training data the samples corresponding with the machine parameters in the $8$ corners of the cube of Fig. 3. This results in $(8\times 10)^{2}=6400$ training samples for each dataset when we pair all curves with each other to form inputs. For testing, we take the remaining $19\times 10$ samples (representing the inner area of the cube and thus allowing a comprehensive analysis) and pair them in turn with each sample corresponding with machine parameters from the $8$ corners, now used as references. This results in $(19\times 10)\times(8\times 10)=15200$ test inputs. The use of training data in testing here does not constitute data leakage, as the references (that have also been used in training data) are an inherent ingredient of the world model as the observed signal that was not included in the training is used to predict the actions in reference to $\mathbf{x}$. We average over the results per individual reference vertex to calculate our evaluation metrics. Additionally, to ensure reliable results, we repeat each experiment $6$ times with different random seeds for model initialization and finally report the mean value derived from these evaluations.
Experiment 2: Sample efficiency. We are interested in learning the model’s performance when only granted minimal training data, considering the time and cost constraints for DOE in the industry. Specifically, we adapt the training setup of Exp. 1 to use only $50$% of the training samples, namely, all samples with machine parameter settings corresponding to the origin and its $3$ directly connected corners (where only one machine parameter varies). This results in $(4\times 10)^{2}=1600$ training inputs when pairing each sample with any other. For evaluation, we take all remaining $23\times 10=230$ samples not used for testing and pair them with the signals resulting from the parameters at the origin as the reference signal, resulting in $(23\times 10)\times(1\times 10)=2300$. Again, we repeat the evaluation $6$ times and report the mean figures of merit.
Experiment 3: Ablation of the latent predictor $P_{z}$. We want to know the effect of the latent predictor $P_{z}$ on the results by comparing the performance of two model variants: one without $P_{z}$ (making the model effectively a direct action regressor) and one with it (identical to the model in the previous experiments). The training and test setup are identical to Exp. 1. This ablation study clarifies whether joint embedding predictive modeling is critical for learning an action-aware representation, leading to improved control.
@@ 5.2 Quantitative evaluation
We aim at evaluating model performance regarding its ability to issue correct control actions, determined by its ability to predict the actions leading from reference signal to observation. This in turn hinges at the model’s ability to accurately disentangle the effects of the individual components (change of one machine parameters) of each action. We quantify this by means of three metrics, detailed below: The action direction $\theta$, action magnitude $d$, and overall prediction quality $q$.
@@
Action direction $\theta$. Does a predicted action point into the correct direction? We consider the ground truth and predicted action vectors $\mathbf{a}$ and $\mathbf{\hat{a}}$, and calculate the angle $\theta(\mathbf{a},\mathbf{\hat{a}})=\arccos\left(\frac{\mathbf{a}\cdot\mathbf{\hat{a}}}{\|\mathbf{a}\|\cdot\|\mathbf{\hat{a}}\|}\right)$ between $\mathbf{a}$ and $\mathbf{\hat{a}}$ (on a scale between 00° to $180$°). A smaller angle indicates better disentanglement, implying that the model can effectively represent the action’s direction in the latent space.
@@
Action magnitude $d$. Does a prediction accurately quantify the magnitude of the needed adjustment? This is answered by the distance between $\mathbf{a}$ and $\mathbf{\hat{a}}$. We compute the absolute distance, where smaller distances indicate more accurate predictions. Specifically, $d(\mathbf{a},\mathbf{\hat{a}})=\sqrt{\sum_{i=1}^{n}(a_{i}-\hat{a}_{i})^{2}}$.
@@
Overall prediction quality $q$. We combine the angle (normalized to [0, 1]) and distance metrics into a single scalar evaluation score to provide a balanced measure of right direction (disentanglement) and magnitude. Specifically, we compute their harmonic mean as $q(\mathbf{a},\mathbf{\hat{a}})=\left(\frac{1}{\theta(\mathbf{a},\mathbf{\hat{a}})}+\frac{1}{d(\mathbf{a},\mathbf{\hat{a}})}\right)^{-1}$. It is well suited as it necessitates the simultaneous minimization of all criteria for good scores.
@@
2D and 3D analysis. We are first interested to see pairwise comparisons between any $2$ action dimensions individually (2D) for fine-grained analysis as this shows how the latent space responds to changes in individual machine parameters. Specifically, for action dimensions $i$ and $j$, let $\mathbf{a_{i,j}}=[a_{i},a_{j}]$, $\mathbf{\hat{a}_{i,j}}=[\hat{a}_{i},\hat{a}_{j}]$, and the 2D angle between them be $\theta(\mathbf{a_{i,j}},\mathbf{\hat{a}_{i,j}})=\arccos\left(\frac{\mathbf{a(i,j)}\cdot\mathbf{\hat{a}(i,j)}}{\|\mathbf{a(i,j)}\|\cdot\|\mathbf{\hat{a}(i,j)}\|}\right)$. Then, we calculate the average angle over all possible 2D pairs as $\theta_{2D}(\mathbf{a},\mathbf{\hat{a}})=\frac{1}{m}\sum_{i,j\in n,i\neq j}\theta(\mathbf{a_{i,j}},\mathbf{\hat{a}_{i,j}})$, where $m$ is the total number of unique pairs (i.e., $m=\frac{n(n-1)}{2}=3$ in our case). The resulting average angle $\theta_{2D}$ serves as a measure of the model’s capacity to disentangle the action dimensions. Correspondingly, 2D distances can be calculated as $d(\mathbf{a_{i,j}},\mathbf{\hat{a}_{i,j}})=\sqrt{\frac{1}{2}[(a_{i}-\hat{a}_{i})^{2}+(a_{j}-\hat{a}_{j})^{2}]}$ and the average distance over all possible 2D pairs as $d_{2D}(\mathbf{a},\mathbf{\hat{a}})=\frac{1}{m}\sum_{i,j\in n,i\neq j}d(\mathbf{a_{i,j}},\mathbf{\hat{a}_{i,j}})$.
@@
All $3$ 2D metrics are used to assess the quality of the predictions regarding individual parameter combinations and the model’s performance in isolating individual parameter changes. Then, the evaluations are extended to all action dimensions simultaneously (3D), analyzing the model’s ability to generalize disentanglement across the full action space. Specifically, using $n=3$ we compute $\theta(\mathbf{a},\mathbf{\hat{a}})$, $d(\mathbf{a},\mathbf{\hat{a}})$ and $q(\mathbf{a},\mathbf{\hat{a}})$ using the formulas from above. This provides insight on how the model handles multi-parameter changes.
6 Results
6.1 Exploring the latent space
@@
Action-awareness in the latent representation. The model resulting according to Exp. 1 on dataset D1 is analyzed regarding properties of its learned latent space. Fig. 4 first shows (all $4$ panels) that the differences between the actual latents ($\boldsymbol{z_{y}}-\boldsymbol{z_{x}}$) are well aligned with the same differences between the reference’s latent and the prediction of the observation’s latent ($\boldsymbol{\hat{z}_{y}}-\boldsymbol{z_{x}}$). This demonstrates the model’s ability to accurately predict activations in latent space based on actions. Moreover, the variation in specific latent dimensions with respect to different actions (all bot top-left panel) highlights the interpretability of the latent space and its sensitivity to action-specific behaviors: It appears action-aware. For example, as depicted in the bottom left corner of Fig. 4, when action is only taken regarding injection speed, the $0^{th}$, $2^{nd}$ and $9^{th}$ dimension change. On the other hand, when action happens regarding mold temperature, only the $1^{st}$ dimension is affected. These effects are mutually exclusive.
@@
Figure 4: Visualization of differences between embeddings (either directly as encoded, blue; or as encoded by $Enc$ for $\mathbf{x}$ and predicted by $P_{z}$ for $\mathbf{y}$, yellow) in the $10$-dimensional embedding space. The red dashed lines indicate a threshold for significantly different activity between the two embeddings. Four different signal pairs are shown: One with a no-op action (top left), and one each for a change only in mold temperature, injection speed or holding pressure (top right, bottom left and bottom right).
Figure 5: Visualization of embeddings from D1’s test set for Exp. 1 in $2$ dimensions using PCA, colored by the actions that led to the underlying signals. Red dots represent embeddings with no action taken, serving as the reference for the other signals; green, blue and purple dots were created by changing only injection speed, molding temperature or holding pressure. The $10$ representations for each of the $4$ actions are clearly separated.
6.2 Action prediction
The evaluation results of Exp. 1–4 on the experiment-specific test sets from datasets D1 and D2 are presented in Tab. 2. The experimental setup is briefly summarized in col. $2$ (with or without $P_{z}$ or transfer learning, number of machine parameter settings in training set). The next $3$ columns give the final metrics for the 2D evaluations, followed by the last $3$ columns for the 3D evaluations.
Table 2: Evaluation of different experimental setups on experiment-specific test sets of D1 and D2
Experiment Experimental setup Avg. action dir. $\mathbf{\theta_{2D}}$ Avg. action mag. $\mathbf{d_{2D}}$ Avg. overall $\mathbf{q_{2D}}$ Avg. action dir. $\mathbf{\theta}$ Avg. action mag. $\mathbf{d}$ Avg. overall $\mathbf{q}$
@@ 6.2 Action prediction
Evaluating Exp. 1–4 on D2. Similar to D1, the model demonstrates poor performance using only $4$ training settings (Exp. 2 vs Exp. 1), as reflected in big angle to and long distance from the ground truth. Increasing the number of settings to $8$ improves performance significantly, showing the importance of geometric context. A slight improvement can be observed when the latent predictor is added (Exp. 3 vs Exp. 1). However, pre-training is harmful no matter if latent predictors are absent or not. For instance, Exp. 4.1 shows a degradation in average $\theta$ from $12.14$° (Exp. 1) to $23.89$° and in average $d$ from $0.20$ to $0.34$. This result is understandable and aligns with the discussion in Sec. 4.2: The limited and noisy pre-training data is from a single source, leading to the learned representation not being generalizable to all cases, specifically when the new case is significantly different from the source domain. Then, it causes negative transfer [2]. The best performance is observed in Exp. 1 with latent predictors and no pre-training, achieving the lowest errors across both 2D ($\theta_{2D}$: $9.59$°, $q_{2D}$: $0.08$) and 3D ($\theta$: $12.14$°, $q$: $0.1$). Overall, combining latent predictors with sufficient training settings consistently improves performance. However, pre-training is not recommended if the pre-training data significantly differs from the data of the target task.
@@
Qualitative evaluation. The model is designed to accurately predict the deviation of machine parameters relative to the reference, regardless of the scale. 2D pairwise qualitative results on D1 from Exp. 1 are depicted in Fig. 6, comparing predicted actions with ground truth. The figures indicate that the effect of actions regarding holding pressure and mold temperature is well isolated, as the predicted actions are close to the ground truth when these actions are taken. However, the effect of injection speed remains entangled with other parameters, indicating that its influence is not as distinctly separable. One reason is that injection speed and holding pressure are highly correlated also in reality, for example because changes in injection speed might influence the material flow which in turn affects pressure.
7 Conclusions
@@
We proposed an actionable world model for industrial process control. It shows potential in disentangled latent representation learning and precise action prediction, improving over the state of the art over prior work [3]. This is achieved by creating a novel methodology that generalizes JEPA to learn action-aware representations for arbitrary, factored actions. The results across multiple datasets highlight the importance of the latent predictor, ensuring better disentanglement and improved prediction quality of actions that can be used as process control signals. Additionally, increasing training data (linearly related to the number of settings) enhances alignment with the ground truth and optimizes model’s prediction quality. However, the results also emphasize necessary caution with respect to pre-training with limited or noisy data. In some cases, it may lead to performance degradation due to negative transfer from data that does not generalize to the case at hand.
@@
Future work. The model has been applied successfully to injection molding cases with $3$ machine parameters. Next, we want to expand this approach to include process datasets covering more machine parameters. Additionally, the effectiveness of transfer learning must be investigated using diverse datasets from different sources, which can be used for downstream fine-tuning. A key challenge is the non-bijective relationship between process signals and machine parameters, as the same process signal may correspond to different parameters. On a larger scale, actionable world models that learn representations of their percepts with respect to issuing goal-driven behavior are a promising avenue for progress in many other fields of AI, from robotics to autonomous agents.

### v2

@@ IV-C Training details
We train the world model using stochastic gradient descent with batch size $32$ for $500$ epochs using the Adam optimizer [35] with learning rate $\alpha=3\times 10^{-4}$ and the default settings otherwise. For fine-tuning, the model is first pre-trained for $10$ epochs. Then, all parameters of $P_{z}$, $P_{a}$ and the last layer of $Enc$ are fine-tuned for $500$ epochs. The coefficients for the loss function in Eq. (4) are set as $\lambda_{1}=1,\lambda_{2}=10$, balancing the terms such that they are of roughly equal magnitude.
V Experimental setup and evaluation
V-A Experiments
@@
Experiment 1: Action space coverage. The goal is to learn to what extent the world model can learn the effects of actions across the whole action space, based on a sparse sampling of all possible actions. We conduct this analysis individually on datasets D1 and D2 that offer a systematic coverage of the action space. We use as training data the samples corresponding with the machine parameters in the $8$ corners of the cube of Fig. 3. Specifically, we pair each of the $8\times 10$ samples with every other, resulting in $6400$ training samples. For testing, we take the remaining $19\times 10$ samples (representing the inner area of the cube and thus allowing a comprehensive analysis) and pair them in turn with each sample corresponding with machine parameters from the $8$ corners as references. This results in $(19\times 10)\times(8\times 10)=15200$ test inputs. The use of training data in testing here does not constitute data leakage, as the references (that are used in training data) are an inherent ingredient of the world model as the observed signal that is not included in the training is used to predict the actions in reference to $\boldsymbol{x}$. We average over the results per individual reference vertex to calculate our evaluation metrics. Additionally, to ensure reliable results, we repeat each experiment $6$ times with different random seeds for model initialization and finally report the mean value derived from these evaluations.
Experiment 2: Sample efficiency. We are interested in learning the model’s performance when only granted minimal training data, considering the time and cost constraints for the DOE in industry. Specifically, we adapt the training setup of Exp. 1 to use only $50$% of the training samples, namely, all samples with machine parameter settings corresponding to the origin and its $3$ directly connected corners (where only one machine parameter varies). This results in $(4\times 10)^{2}=1600$ training inputs when pairing each sample with any other. For evaluation, we take all remaining $23\times 10=230$ samples not used for testing and pair them with the signals resulting from the parameters at the origin as the reference signal, resulting in $(23\times 10)\times(1\times 10)=2300$. Again, we repeat the evaluation $6$ times and report the mean figures of merit.
Experiment 3: Ablation of the latent predictor $P_{z}$. We assess the effect of the latent predictor $P_{z}$ on the results by comparing the performance of two model variants: one without $P_{z}$ (making the model effectively a direct action regressor) and one with it (identical to the model in the previous experiments). The training and test setup are identical to Exp. 1. This ablation study clarifies whether joint embedding predictive modeling is critical for learning an action-aware representation, leading to improved control.
@@ V-B Quantitative evaluation
We aim at evaluating model performance regarding its ability to issue correct control actions, determined by its ability to predict the actions leading from reference signal to observation. This in turn hinges on the model’s ability to accurately disentangle the effects of the individual components (change of one machine parameters) of each action. We quantify this by means of three metrics, detailed below: The action direction $\theta$, action magnitude $d$, and overall prediction quality $q$.
@@
Action direction $\theta$. Does a predicted action point into the correct direction? We consider the ground truth and predicted action vectors $\boldsymbol{a}$ and $\boldsymbol{\hat{a}}$, and calculate the angle $\theta(\boldsymbol{a},\boldsymbol{\hat{a}})=\arccos\left(\frac{\boldsymbol{a}\cdot\boldsymbol{\hat{a}}}{\|\boldsymbol{a}\|\cdot\|\boldsymbol{\hat{a}}\|}\right)$ between $\boldsymbol{a}$ and $\boldsymbol{\hat{a}}$ (on a scale between 00° to $180$°). A smaller angle indicates better disentanglement, implying that the model can effectively represent the action’s direction in the latent space.
@@
Action magnitude $d$. Does a prediction accurately quantify the magnitude of the needed adjustment? This is answered by the distance between $\boldsymbol{a}$ and $\boldsymbol{\hat{a}}$. We compute the absolute distance, where smaller distances indicate more accurate predictions. Specifically, $d(\boldsymbol{a},\boldsymbol{\hat{a}})=\sqrt{\sum_{i=1}^{n}(a_{i}-\hat{a}_{i})^{2}}$, where $n$ denotes the dimension of the vectors.
@@
Overall prediction quality $q$. We combine the angle (normalized to [0, 1]) and distance metrics into a single scalar evaluation score to provide a balanced measure of right direction (disentanglement) and magnitude. Specifically, we compute their harmonic mean as $q(\boldsymbol{a},\boldsymbol{\hat{a}})=\frac{1}{2}\left(\frac{1}{\theta(\boldsymbol{a},\boldsymbol{\hat{a}})}+\frac{1}{d(\boldsymbol{a},\boldsymbol{\hat{a}})}\right)^{-1}$. It is well suited as it necessitates the simultaneous minimization of all criteria for good scores.
@@
2D and 3D analysis. Our evaluation employs complementary 2D and 3D analyses to examine the model’s action prediction capabilities at different levels of complexity. We are first interested to see pairwise comparisons between any $2$ action dimensions individually (2D) for fine-grained analysis as this shows how the latent representation responds to changes in individual machine parameters. Specifically, for action dimensions $i$ and $j$, let $\boldsymbol{a_{i,j}}=[a_{i},a_{j}]$, $\boldsymbol{\hat{a}_{i,j}}=[\hat{a}_{i},\hat{a}_{j}]$, and the 2D angle between them be $\theta(\boldsymbol{a_{i,j}},\boldsymbol{\hat{a}_{i,j}})=\arccos\left(\frac{\boldsymbol{a(i,j)}\cdot\boldsymbol{\hat{a}(i,j)}}{\|\boldsymbol{a(i,j)}\|\cdot\|\boldsymbol{\hat{a}(i,j)}\|}\right)$. Then, we calculate the average angle over all possible 2D pairs as $\theta_{2D}(\boldsymbol{a},\boldsymbol{\hat{a}})=\frac{1}{m}\sum_{i,j\in n,i\neq j}\theta(\boldsymbol{a_{i,j}},\boldsymbol{\hat{a}_{i,j}})$, where $m$ is the total number of unique pairs (i.e., $m=\frac{n(n-1)}{2}=3$ in our case). The resulting average angle $\theta_{2D}$ serves as a measure of the model’s capacity to disentangle the action dimensions. Correspondingly, 2D distances can be calculated as $d(\boldsymbol{a_{i,j}},\boldsymbol{\hat{a}_{i,j}})=\sqrt{\frac{1}{2}[(a_{i}-\hat{a}_{i})^{2}+(a_{j}-\hat{a}_{j})^{2}]}$ and the average distance over all possible 2D pairs as $d_{2D}(\boldsymbol{a},\boldsymbol{\hat{a}})=\frac{1}{m}\sum_{i,j\in n,i\neq j}d(\boldsymbol{a_{i,j}},\boldsymbol{\hat{a}_{i,j}})$.
@@
All $3$ 2D metrics assess the prediction quality of the individual parameter combinations and the model’s performance in isolating individual parameter changes. Then, the evaluations are extended to all action dimensions simultaneously (3D), analyzing the model’s ability to generalize disentanglement across the full action space. Specifically, we compute $\theta(\boldsymbol{a},\boldsymbol{\hat{a}})$, $d(\boldsymbol{a},\boldsymbol{\hat{a}})$ and $q(\boldsymbol{a},\boldsymbol{\hat{a}})$ by extending the vectors in 3D using the formulas from above. This provides insight on how the model handles multi-parameter changes.
VI Results
VI-A Exploring the latent space
@@
Action-awareness in the latent representation. The model resulting from Exp. 1 on dataset D1 is analyzed regarding properties of its learned latent space. Fig. 4 first shows (all $4$ panels) that the differences between the actual latents ($\boldsymbol{z_{y}}-\boldsymbol{z_{x}}$) are well aligned with the same differences between the reference’s latent and the prediction of the observation’s latent ($\boldsymbol{\hat{z}_{y}}-\boldsymbol{z_{x}}$). This demonstrates the model’s ability to accurately predict activations in latent space based on actions, enabling simulation of the latent space for arbitrary actions. Moreover, the variation in specific latent dimensions with respect to different actions (all but top-left panel) highlights the interpretability of the latent space and its sensitivity to action-specific behaviors: It appears action-aware. For example, as depicted in the bottom left corner of Fig. 4, when action is only taken regarding injection speed, the $0^{th}$, $2^{nd}$ and $9^{th}$ dimension change. On the other hand, when the mold temperature is adjusted, only the $1^{st}$ dimension is significantly affected. These effects are mutually exclusive.
@@
Figure 4: Visualization of differences between embeddings (either directly as encoded, blue; or as encoded by $Enc$ for $\boldsymbol{x}$ and predicted by $P_{z}$ for $\boldsymbol{y}$, orange) in the $10$-dimensional embedding space. The red dashed lines indicate a threshold for significantly different activity between the two embeddings. Four different signal pairs are shown: One with a no action (top left), and one each for a change only in mold temperature, injection speed or holding pressure (top right, bottom left and bottom right).
Figure 5: Visualization of embeddings from D1’s test set for Exp. 1 in $2$ dimensions using PCA, colored by the actions that led to the underlying signals. Red dots represent embeddings with no action taken, serving as the reference for the other signals; green, blue and purple dots are created by changing only injection speed, molding temperature or holding pressure.
VI-B Action prediction
The evaluation results of Exp. 1–4 on the experiment-specific test sets from datasets D1 and D2 are presented in Tab. II. The experimental setup is briefly summarized in col. $2$ (with or without $P_{z}$ or transfer learning, number of machine parameter settings in training set). The next $3$ columns give the final metrics for the 2D evaluations, followed by the last $3$ columns for the 3D evaluations.
TABLE II: Evaluation of different experimental setups on experiment-specific test sets of D1 and D2
Experiment Experimental setup Avg. action dir. $\boldsymbol{\theta_{2D}}$ Avg. action mag. $\boldsymbol{d_{2D}}$ Avg. overall $\boldsymbol{q_{2D}}$ Avg. action dir. $\boldsymbol{\theta}$ Avg. action mag. $\boldsymbol{d}$ Avg. overall $\boldsymbol{q}$
@@ VI-B Action prediction
Evaluating Exp. 1–4 on D2. Similar to D1, the model demonstrates poor performance using only $4$ training settings (Exp. 2 vs Exp. 1), as reflected in big angle to and long distance from the ground truth. Increasing the number of settings to $8$ improves performance significantly, showing the importance of geometric context. A slight improvement can be observed when the latent predictor is added (Exp. 3 vs Exp. 1). However, pre-training is harmful no matter if latent predictors are absent or not. For instance, Exp. 4.1 shows a degradation in average $\theta$ from $12.14$° (Exp. 1) to $23.89$° and in average $d$ from $0.20$ to $0.34$. This result is understandable and aligns with the discussion in Sec. IV-B: The limited and noisy pre-training data is from a single source, leading to the learned representation not being generalizable to all cases, specifically when the new case is significantly different from the source domain. Then, it causes negative transfer [2]. The best performance is observed in Exp. 1 with latent predictors and no pre-training, achieving the lowest errors across both 2D ($\theta_{2D}$: $9.59$°, $q_{2D}$: $0.08$) and 3D ($\theta$: $12.14$°, $q$: $0.1$). Overall, combining latent predictors with sufficient training settings consistently improves performance. However, pre-training is not recommended if the pre-training data significantly differs from the data of the target task.
@@
Qualitative evaluation. The model is designed to accurately predict the deviation of machine parameters relative to the reference, regardless of the scale. 2D pairwise qualitative results on D1 from Exp. 1 are depicted in Fig. 6, comparing predicted actions with ground truth. The figures indicate that the effect of actions regarding holding pressure and mold temperature is well isolated, as the predicted actions are close to the ground truth when these actions are taken. However, the effect of injection speed remains entangled with other parameters, indicating that its influence is not as distinctly separable. A key reason is that injection speed and holding pressure are highly correlated in reality, for example, changes in injection speed might influence the material flow which in turn affects pressure.
VII Conclusions
@@
We proposed an actionable world model for industrial process control. It shows potential in disentangled latent representation learning and precise action prediction, improving over the state of the art over prior work [3]. This is achieved by creating a novel methodology that generalizes JEPA to learn action-aware representations for arbitrary, factored actions. The results across multiple datasets highlight the importance of the latent predictor, ensuring better disentanglement and improved prediction quality of actions that can be used as process control signals. Additionally, increasing training data (linearly related to the number of settings) enhances alignment with the ground truth and optimizes model’s prediction quality. However, the results also emphasize necessary caution with respect to pre-training with limited or unrelated data. In some cases, it may lead to performance degradation due to negative transfer from data that do not generalize to the case at hand.
@@
Future work. The model has been applied successfully to injection molding cases with $3$ machine parameters. Next, we want to expand this approach to include process datasets covering more machine parameters. Furthermore, it would be also interesting to expand this framework for other industrial applications. Additionally, the effectiveness of transfer learning must be investigated using diverse datasets from different sources, which can be used for downstream fine-tuning. A key challenge is the non-bijective relationship between process signals and machine parameters, as the same process signal may correspond to different parameters. On a larger scale, actionable world models that learn representations of their percepts with respect to issuing goal-driven behavior are a promising avenue for progress in many other fields of AI, from robotics to autonomous agents.